# Divide-Then-Aggregate: An Efficient Tool Learning Method via Parallel Tool Invocation

### v1

@@ 4 Experimental Setup
We use StableToolBench Guo et al. (2024) for evaluation. All test cases in StableToolBench are actually derived from the test portion of ToolBench Touvron et al. (2023b). Concretely, ToolBench is divided into six evaluation subsets based on tool categories and scenarios. The tool categories are as follows: Inst. denotes unseen instructions for the same set of tools in the training data, Tool denotes unseen tools within the same (seen) category as those in the training data, and Cat. denotes unseen tools from a different (unseen) category. The scenarios are: I1 for single-tool instructions, I2 for intra-category multi-tool instructions, and I3 for intra-collection multi-tool instructions. The difficulty level of the task escalates progressively from I1-Inst. to I3-Inst.. Comparing to the original ToolBench, StableToolBench introduces an extra caching system and an API simulator to mitigate the instability issues associated with real-time APIs.
@@
[…]
@@
We use the Llama-series models Touvron et al. (2023a, b); Dubey et al. (2024) as our backbone, fine-tuning them for our task. Since Llama2 has a context length limit of 4096 tokens, shorter contexts may not be sufficient for effective tool invocation. To overcome this limitation, we followed the approach in ToolLLaMA Qin et al. (2023), and applied position interpolation Chen et al. (2023) to extend the context length to 8192 tokens. More details about out training process can be found in Appendix D.
@@ 5.1 Main Experiments
Furthermore, under the same GPT model conditions, DFSDT and Parallel show no significant performance differences, indicating that their parallel function call strategy does not significantly improve the tool invocation capabilities of GPT LLMs. In the open-source LLMs, Qwen2.5 (Parallel) demonstrated a clear improvement over DFSDT, and our method further advances this performance. This confirms that our parallel mechanism is more effective than other parallel approaches, offering a greater boost to tool invocation capabilities.
@@
The Solvable Pass Rate evaluates the quality of results against the GPT-3.5 (ReAct) baseline. As shown in Table 2, most open-source models achieve a SoWR score below 50, indicating their responses are of lower quality compared to GPT-3.5 (ReAct). In contrast, our method outperforms all open-source models, achieving a SoWR score of 70.8 on the I2-Inst dataset (27.32% relative improvement). These results highlight the effectiveness of our framework. Further analysis reveals that this improvement can be attributed to our parallel tool invocation mechanism, which broadens the scope of LLMs, thereby shortening the overall trajectory and reducing the input load, resulting in better responses.
Compared to GPT-3.5, at around 7B parameters, our method provides performance that is nearly equivalent, and in some cases, even exceeds it. However, we acknowledge a performance gap when compared to GPT-4’s official parallel method. A possible explanation lies in the foundational capabilities of our backbone, LLama2-7B, which currently falls short of GPT-4 in areas such as long-term planning and long-context reasoning.
@@ 5.2 Computational Cost
We count the tokens for both Completion and Prompt. Completion refers to the output tokens generated by LLMs, while Prompt refers to the input tokens provided to the LLMs. Typically, the cost of Completion is higher than that of Prompt. Figure 3 shows the token consumption for all methods. For each subset of StableToolBench, we calculate the average token count across all cases to assess the overall performance of the LLMs. The results in Figure 3 demonstrate that our approach is highly competitive among open-source models. It substantially outperforms DFSDT while costing less than ReAct. Compared to GPT-based methods, our approach is particularly cost-effective. It consumes fewer tokens, and given that DTA-Llama is a 7-billion-parameter model, the actual deployment costs will be even lower. In addition, we also provide statistics on the maximum token consumption for each subset in Appendix F, to assess the performance of the LLMs in handling complex scenarios.
Furthermore, we analyze the inference step of each baselines. Specifically, we calculate the average number of inference steps for each baselines across all subsets. As demonstratd in Table 3, our method maintains its strong performance in terms of token consumption, requiring the fewest inference steps. In contrast, both the GPT series and Qwen2.5 models exhibit higher inference steps with their Parallel methods, which may be attributed to their limited task-planning capabilities.
@@
[…]
@@
I1-Inst. 34 76↑124% 54↑59% 113↑232% 104↑205% 29↓15%
I1-Tool 40 103↑158% 57↑43% 128↑220% 111↑177% 33↓20%
I1-Cat. 35 90↑157% 94↑169% 146↑317% 84↑140% 28↓20%
I2-Inst. 39 110↑182% 56↑44% 102↑162% 153↑292% 29↓26%
I2-Cat. 38 124↑226% 52↑37% 140↑268% 130↑241% 30↓21%
I3-Inst. 46 120↑161% 81↑76% 139↑202% 135↑193% 40↓13%
Avg. 39 104↑167% 66↑69% 128↑228% 119↑204% 31↓21%
@@
I1-Inst. 1.00 $\times$0.45 $\times$0.63 $\times$0.30 $\times$0.33 $\times$1.18
I1-Tool 1.00 $\times$0.39 $\times$0.37 $\times$0.24 $\times$0.36 $\times$1.20
I1-Cat. 1.00 $\times$0.39 $\times$0.70 $\times$0.31 $\times$0.42 $\times$1.20
I2-Inst. 1.00 $\times$0.31 $\times$0.73 $\times$0.27 $\times$0.26 $\times$1.27
I2-Cat. 1.00 $\times$0.35 $\times$0.69 $\times$0.38 $\times$0.29 $\times$1.35
I3-Inst. 1.00 $\times$0.38 $\times$0.57 $\times$0.33 $\times$0.34 $\times$1.15
Avg. 1.00 $\times$0.37 $\times$0.59 $\times$0.30 $\times$0.33 $\times$1.22
@@ Appendix A Prompt
Figure 5 shows the prompts used during the data transformation process. We use this prompt with GPT-4-turbo to transform the serial data into DAG structure. The prompt first evaluates the reasonableness of the conversation, discarding any data that doesn’t meet the criteria. It then assesses whether the steps can be processed in parallel, creating a planning path in DAG structure. The evaluations of reasonableness and parallelism follow distinct analytical processes, embodying a form of internalized CoT prompt engineering. Applying these analyses to DAG generation leads to more accurate results.
@@
Figure 6 illustrates the system prompt used during training and inference. In ToolBench, each user instruction is paired with a list of tool candidates, comprising both relevant and unrelated tools. The tools includes the names and descriptions of all tools in the candidate set. Additionally, each tool is equipped with a set of APIs designed to handle different types of tasks. The API is represented as a JSON list containing detailed information about the names, descriptions, and parameters of the APIs associated with each tool.
@@ Appendix E Case Study
In Figure 9, we present several cases to demonstrate the performance of DTA-Llama on practical tasks. All the tasks and tools in the figure are sourced from ToolBench. The figure presents two cases, each consisting of the user’s Question (instruction), the LLMs’ Output and the Tool Response. In DTA-Llama, if necessary, the system concurrently executes multiple tool strategies during the LLM Output phase and aggregates the results of all tool invocation during the Tool Response phase.
@@ Appendix F Supplementary of the Experiments
In Table 5, we present the maximum token consumption for each method across various subsets. As shown in the table, the ReAct method has the lowest maximum token count, with our method following closely behind. The ReAct method employs a simple and straightforward task-planning mechanism, resulting in low resource consumption but the lowest success rate. It frequently struggles with complex tasks. In contrast, our method efficiently leverages additional tokens to tackle more challenging tasks without substantially increasing token usage, demonstrating superior cost-effectiveness.

### v2

@@ 4 Experimental Setup
We use StableToolBench Guo et al. (2024) for evaluation. All test cases in StableToolBench are actually derived from the test portion of ToolBench Touvron et al. (2023b). Concretely, ToolBench is divided into six evaluation subsets based on tool categories and scenarios. The tool categories are as follows: Inst. denotes unseen instructions for the same set of tools in the training data, Tool denotes unseen tools within the same (seen) category as those in the training data, and Cat. denotes unseen tools from a different (unseen) category. The scenarios are: I1 for single-tool instructions, I2 for intra-category multi-tool instructions, and I3 for intra-collection multi-tool instructions. The difficulty level of the task escalates progressively from I1-Inst. to I3-Inst.. Compared to the original ToolBench, StableToolBench introduces an extra caching system and an API simulator to mitigate the instability issues associated with real-time APIs.
@@
We use both GPT-series and other open-source LLMs as our baselines. For the GPT-series models, we use OpenAI’s GPT-3.5-turbo and GPT-4-turbo, leveraging their function calling capabilities***While the exact mechanisms remain unclear, OpenAI has enabled parallel tool invocation in these models.. And we include Parallel as an additional baseline paradigm alongside ReAct and DFSDT.
For open-source models, we fine-tune ToolLLaMA (Qin et al., 2023) from Llama2-7B on ToolBench and compare it using the ReAct and DFSDT methods. LLMCompiler Kim et al. (2023) is a non-training-based parallel tool invocation method that relies on system design and prompt engineering. We also use it as a baseline for comparison with Llama2-7B. Additionally, recent open-source LLMs, such as Qwen2.5 Qwen Team (2024), have demonstrated strong capabilities, including function calls and parallel tool invocation. Consequently, we include Qwen2.5-7B-Instruct in our baseline for comparison. To ensure experimental consistency, our method is fine-tuned on Llama2-7B for a fair comparison.
@@
We use the Llama-series models Touvron et al. (2023a, b); Dubey et al. (2024) as our backbone, and fine-tune them for our task. Since Llama2 has a context length limit of 4096 tokens, shorter contexts may not be sufficient for effective tool invocation. To overcome this limitation, we followed the approach in ToolLLaMA Qin et al. (2023) and applied position interpolation Chen et al. (2023) to extend the context length to 8192 tokens. More details about our training process can be found in Appendix D.
@@ 5.1 Main Experiments
Moreover, under the same GPT model conditions, DFSDT and Parallel demonstrate similar performances, suggesting that their parallel function call strategy does not notably enhance the tool invocation capabilities of GPT-based LLMs. In contrast, Qwen2.5 (Parallel) exhibited a clear improvement over DFSDT in open-source LLMs, with our method further advancing this performance. This demonstrates that our parallel mechanism is more effective than other parallel approaches, providing a greater boost to tool invocation capabilities.
@@
The Solvable Pass Rate evaluates the quality of results against the GPT-3.5 (ReAct) baseline. As shown in Table 2, most open-source models achieve a SoWR score below 50, indicating their responses are of lower quality compared to GPT-3.5 (ReAct). In contrast, our method outperforms all open-source models, achieving a SoWR score of 70.8 on the I2-Inst dataset (27.32% relative improvement). Compared to GPT-3.5 DFSDT & Parallel, our method achieves nearly equivalent performance based solely on Llama2-7B, and even surpasses it in certain subsets.
These results, combined with SoPR, demonstrate that our method significantly enhances the ability of LLMs’ tool utilization. We attribute this improvement to our specialized parallel tool invocation mechanism. This mechanism expands the decision-making scope of LLMs within one round, reducing the overall trajectories while enabling better decision-making. We provide example of the actual decision-making processes of different methods in the Appendix E to illustrate this.
@@ 5.2 Computational Cost
We count tokens for both Completion (output tokens generated by LLMs) and Prompt (input tokens provided to LLMs). Typically, Completion is more costly than Prompt. Figure 3 shows the token consumption for all methods. For each subset of StableToolBench, we calculate the average token count across all cases to assess the overall performance of the LLMs. The results in Figure 3 show that our approach is highly competitive among open-source models. It significantly outperforms DFSDT while costing less than ReAct. Compared to GPT-based methods, our approach is particularly cost-effective, consuming fewer tokens. Given that DTA-Llama is a 7-billion-parameter model, its actual deployment costs are even lower. In addition, we also provide statistics on the maximum token consumption for each subset in Appendix F, to assess the performance of the LLMs in handling complex scenarios.
Furthermore, we analyze the inference steps of each baseline by calculating the average number of steps across all subsets. As demonstratd in Table 3, our method maintains its strong performance in terms of token consumption, requiring the fewest inference steps. In contrast, both the GPT series and Qwen2.5 models exhibit higher inference steps with their Parallel methods, which may be attributed to their limited task-planning capabilities. The LLMCompiler, a non-training approach, faces a bottleneck in processing complex tasks with 7-billion-parameter models. When considered alongside the results presented in Table 2, its performance falls short of expectations.
@@
Dataset ReAct DFSDT LLMCompiler Qwen2.5 Ours
@@
I1-Inst. 34 76↑124% 58↑71% 104↑205% 29↓15%
I1-Tool 40 103↑158% 67↑68% 111↑177% 33↓20%
I1-Cat. 35 90↑157% 61↑74% 84↑140% 28↓20%
I2-Inst. 39 110↑182% 83↑113% 153↑292% 29↓26%
I2-Cat. 38 124↑226% 71↑87% 130↑241% 30↓21%
I3-Inst. 46 120↑161% 69↑50% 135↑193% 40↓13%
Avg. 39 104↑167% 68↑74% 119↑204% 31↓21%
@@
I1-Inst. 1.00 $\times$0.45 $\times$0.59 $\times$0.33 $\times$1.18
I1-Tool 1.00 $\times$0.39 $\times$0.60 $\times$0.36 $\times$1.20
I1-Cat. 1.00 $\times$0.39 $\times$0.57 $\times$0.42 $\times$1.20
I2-Inst. 1.00 $\times$0.31 $\times$0.47 $\times$0.26 $\times$1.27
I2-Cat. 1.00 $\times$0.35 $\times$0.54 $\times$0.29 $\times$1.35
I3-Inst. 1.00 $\times$0.38 $\times$0.67 $\times$0.34 $\times$1.15
Avg. 1.00 $\times$0.37 $\times$0.57 $\times$0.33 $\times$1.22
@@ Appendix A Prompt
Figure 5 shows the prompts used during the data transformation process. We use this prompt with GPT-4-turbo to convert the serial data into a DAG structure. The prompt first evaluates the reasonableness of the conversation, discarding any data that doesn’t meet the criteria. It then assesses whether the steps can be processed in parallel, generating a planning path within the DAG structure. The evaluations of reasonableness and parallelism follow distinct analytical processes, reflecting a form of internalized CoT prompt engineering. By applying these analyses to DAG generation, we achieve more accurate results.
@@
Figure 6 illustrates the system prompt used during both training and inference. Each user instruction is paired with a list of tool candidates, including both relevant and unrelated. The tools section contains the names and descriptions of all tools in the candidate set. Additionally, each tool is equipped with a set of APIs designed to handle various types of tasks. The API section is represented as a JSON list, providing detailed information about the names, descriptions, and parameters of the APIs associated with each tool.
@@ Appendix C DTA-Tool Instance
Figure 9: A comparison of the real decision-making visualization processes between CoT/React, DFSDT, and our method: (a) The serial approach is straightforward but prone to getting stuck in local loops when errors occur, leading to task failure; (b) The tree-based decision-making approach attempts various possibilities through deep traversal and backtracking, but due to its narrow decision scope, it involves a large amount of redundant processes; (c) Our method maximizes the parallel use of available tools during the Threads phase, enhancing the perception of LLMs and resulting in an efficient and accurate final outcome.
@@ Appendix E Case Study
Comparison of decision process
Figure 9 illustrates the actual tool invocation decision-making processes of ReAct, DFSDT, and our method.
Complete DTA-Llama performance
In Figure 10, we present several cases to showcase the performance of DTA-Llama on practical tasks. Each case consists of the user’s Question (instruction), the LLM’s Output, and the Tool Response. The LLM’s Output represents the Process, which includes the LLMs’ thought and tool invocation strategy. After the execution of Threads, the Tool Response presents the results of the tools’ execution.
@@ Appendix F Supplementary of the Experiments
In Table 5, we present the maximum token consumption for each method across various subsets. As shown, the ReAct method has the lowest maximum token count, with our method following closely behind. The ReAct method employs a simple and straightforward task-planning mechanism, leading to low resource consumption but a very low success rate, which makes it ill-suited for handling complex real-time tasks. In contrast, our method efficiently utilizes additional tokens to tackle more challenging tasks without significantly increasing token usage, demonstrating superior cost-effectiveness.